OpenStudy (anonymous):

Is y=x^2 a function? Is the inverse of y=x^2 a function? Justify your answer

OpenStudy (anonymous):

Do you know the inverse of y=x^2 first of all?

OpenStudy (anonymous):

Is it x = y?

OpenStudy (anonymous):

swap the x and y in f(x) to get f^-1(x) x=y^2, solve for y

OpenStudy (anonymous):

\[y= \pm \sqrt{x}\]

OpenStudy (anonymous):

so you're going to have 2 y values for each x-value that is not 0

OpenStudy (anonymous):

is that a function?

OpenStudy (anonymous):

In a function, each member of the domain corresponds to at most 1 member of the range

OpenStudy (anonymous):

No it's not.

OpenStudy (anonymous):

right, and is y=x^2 a function?

OpenStudy (anonymous):

No, because it's a parabola?

OpenStudy (anonymous):

it is a parabola, as is the inverse, but in this case, each x in the domain (all real numbers) corresponds to exactly one member of the range

OpenStudy (anonymous):

You can use the vertical line test

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):